OpenStudy (newarf):

add the fractions and write your answer in the lowest terms 2/5 + 1/9

563blackghost (563blackghost):

First convert these two fraction to be with the same denominator....we can multiply each equation with the opposite denominator to have them the same... \(\huge{\frac{2}{5} \times \frac{9}{9}=?}\) \(\huge{\frac{1}{9} \times \frac{5}{5}=?}\)